# ManiBox: Enhancing Spatial Grasping Generalization via Scalable Simulation Data Generation

### v1

@@ Appendix B Ablation Study
Appendix C Teacher-Student Training Details
C.1 Teacher Policy Details
@@
We train our teacher policy using PPO with 8,192 parallel environments in Isaac Lab (Mittal et al., 2023). With a carefully designed reward function and applied PPO techniques, the training is completed in around 40 minutes, reaching approximately 2,000 training iterations.
C.1.1 Simulation Scenario Setup
@@
The environment is built based on the lift task in Isaac Lab. During training, the scene consists of a table, an apple, and a robot. To support future extensions to more complex tasks, such as obstacle avoidance during grasping, we represent the robot’s behavior with a 30Hz trajectory. Each environment in the simulator is evenly spaced to prevent interference between different training instances. The apple’s scale is uniformly randomized across environments, while its spatial position and the height of the table are uniformly randomized for each individual run. Further details on the randomized ranges can be found in Table 3.
@@
C.1.2 Details of Observation and Action
@@ C.1.3 Implementation of PPO
C.1.4 Reward Function Design
To ensure the policy achieves the desired grasping behavior on our robot, we design a reward function that reflects the task’s objectives. As grasping is complicated, it can be divided into a multi-stage reward function: first reaching the object, then closing the jaws, then picking up the object, and finally retracting the robotic arm to a specified position.
Details of our reward function design for the specific object-grasp tasks can be found in Table 5
@@ C.1.4 Reward Function Design
C.2 Data Generation Details
@@
To reduce visual interference between parallel environments and minimize the sim-to-real gap, we incorporate a white room during data collection. Three RGB-D cameras [^1] But we only use RGB information., calibrated with intrinsic and extrinsic parameters matching those of the real robot, capture visual information. During inference, the images from these cameras are processed by YOLO-World to generate object bounding boxes.
@@
C.3 Student Policy Details
@@
Appendix D Hardware Details
The robot is equipped with a dual-arm design, providing a total of 14 degrees of freedom (DoF), with 7 DoF per arm for versatile movement. Each arm has a maximum valid payload of 1500g, and the gripper has a range of up to 80mm, allowing the robot to perform a variety of grasping and manipulation tasks efficiently. The detailed hardware information and an image of the robot can be found in Figure 8 and Table 10.
@@ Appendix D Hardware Details
Appendix E Experimental Details
E.1 Evaluation in Simulation
We run large-scale experiments in simulator to test the spatial generalization curve for each range and determine the minimum data amounts to achieve optimal success rate. To align with real machine inference, we set robot’s joints directly to target qpos instead of using environment stepping. We mark the trajectory as success if there is a step both reaching reward $>0.6$ and quat reward $>1.4$. For each policy, we test 48 trajectories and record the final success rate.
E.2 Experiment Result Analysis
We discuss here the analysis of the results of spatial scaling law.
[…]
For the functional relationship between the amount of data required for spatial generalization and spatial volume, we explored two types of fits: a power law and a logarithmic function. The fitting results of power law is given by $y=a\cdot x^{b}=1507x^{0.28}$, while the logarithmic fit is $y=a\cdot\ln(b\cdot x)+c=2450.69\cdot\ln{(0.51\cdot x)}-1604.28$. We found that the logarithmic fit becomes unreasonable as the spatial volume $x$ approaches 0, where the required data, $y=2450.69\cdot\ln{0.51\cdot x}-1604.28$, tends to negative infinity, which is unrealistic since the amount of data cannot be negative. Furthermore, the fit results in a negative value at $x=2$, where $bx=1$ and $\log(bx)=0$, leading to a required data amount of $a\log(bx)+c=-1604$. In the long run (for larger spatial volumes), the power law fit better describes the relationship between the data required for spatial generalization and spatial volume. Fig. 9 shows the selected power law relationship, and Fig. 10 displays the logarithmic curve.
E.3 Evaluation in Real World
To verify our experiment results from simulator, we also test our policy in real robot. For each range, we choose the first optimal policy that achieves 80% success rate in simulator and evaluate with real machine for 10 randomly chosen points.
In addition to the verification of spatial generalization, we also test the ability of object, background, and complex generalizations. The details of experiment setting can be found in Table 11.
Table 11: Real-World Experiment Settings
Generalization Type Setting Description
Object Generalization
Yellow Pear h4: Point Random, Minimal setting
Orange h4: Point Random, Minimal setting
Steel Cup h4: Point Random, Minimal setting
Tin Can h4: Point Random, Minimal setting
Glass Beaker h4: Point Random, Minimal setting
Background Generalization
Apple h4: Point Random, Orange tablecloth
Unseen Apple h4: Point Random, Black tablecloth
Apple h4: Point Random, Playing video
Apple h4: Point Random, Dim light
Apple h4: Point Random, Flittering light
Complex Generalization
Unseen Apple, Unseen Table Point Random, Table 2
Unseen Apple, Unseen Desk Point Random, Multiple objects
Steel Cup, Unseen Desk Point Random, Daily desktop
Plastic Bottle, Unseen Desk Point Random, Daily desktop
The three camera images, as well as the entire trajectory of grasping, are schematized below:

### v2

@@ Appendix B Ablation Study
B.5 Comparison with IK and Other Baselines in the Simulator
In order to demonstrate the robustness of our method compared to other baseline methods on the grasping task, we compare our method to a heuristics-based approach, in particular the Inverse Kinematics (IK) algorithm from Isaac Lab which is required in some heuristics-based approaches such as AnyGrasp (Fang et al., 2023). We conduct tests using 3 different seeds in the simulator, and the comparison results are as follows:
  • Inverse Kinematics (IK): This approach directly uses the ground truth object pose as the target end-effector (eef) pose, then computes the joint positions via IK to move the robot arm to the target. However, we observe that IK performs poorly at the edges of the workspace, especially when objects are near the table surface or at low positions. Examples of IK failures when the table is very low can be seen in Fig. 10 and Fig. 11. An example of an IK success can be seen in Fig. 12. The reason for IK failures is possible:
    • The robotic arm may approach singular configurations where the Jacobian matrix becomes singular, causing unstable or undefined IK solutions.
    • Near the workspace edge, numerical inaccuracies may result in imprecise IK solutions. Moreover, approximation methods (e.g., pseudoinverse) may struggle to converge to a valid solution in complex boundary cases.
  • Key Steps: In this approach, only key steps of the trajectory are used to model the grasping task.
  • Without Random Mask: This variant trains the student policy without random masking for object positioning.
Table 4: Success Rate of More Baseline Methods on the Full Space of the Simulator.
Method Success Rate of Full Space
Inverse Kinematics (IK) $68.75\%\pm 5.10\%$
Key Steps $6.25\%\pm 8.84\%$
Without Random Mask $58.33\%\pm 7.80\%$
Student Policy of ManiBox (Ours) $91.67\%\pm 2.95\%$
Here we get some key insights:
  • IK Limitations: As expected, IK struggles in the workspace boundaries due to singularities, which cause failures when the object is positioned at the edges of the table or at low heights. This limitation is inherent to traditional geometric methods like IK, which rely on precise pose knowledge and are prone to errors in challenging environments.
  • Our RL-Based Approach: In contrast, ManiBox (Ours) utilizes RL to learn robust grasping teacher policies by directly planning the joint positions, enabling successful manipulation across a wide range of object positions. Our method does not rely on explicit pose knowledge and is highly effective even at challenging object locations that cause IK to fail.
Figure 10: An example of IK failure (view of the center camera): the object is at the edge of the robotic arm’s workspace - a case where the object height is too low.
Figure 11: An example of IK failure (view of the left camera): the object is at the edge of the robotic arm’s workspace - a case where the object height is too low.
Figure 12: An example of an IK success.
Appendix C More Real-World Experiments
To demonstrate the extensibility of our approach, we supplement four more real-world experiments: Pour Water (Figure 13), showcasing our approach’s extensibility to multi-object manipulation tasks; Grab the handle of the cup (Figure 14), illustrating our approach’s extensibility to the grasping of detailed parts of irregular objects; Mid-Air (Figure 15); and Cluttered Table at varying heights (Figure 16 and Figure 17).
Figure 13: Task: Pour Water. The policy grasps the left water bottle and pours water into the right cup, demonstrating the extensibility of our approach to complex tasks. By modifying the teacher policy in the simulator, this method can be extended to more sophisticated tasks with more objects.
Figure 14: Task: Grab the handle of the cup. More fine-grained detection can be achieved using models such as Liu et al. (2023), so that our policy can catch irregular objects, such as the handle of a cup. The first figure shows the recognition of the cup handle from the first-person view and the other figures show the task completion process from the third-person view.
Figure 15: Task: Mid-Air. Grasp the bottle lifted by hand and suspended in mid-air, then move it to the target position.
Figure 16: Task: Cluttered Table. Grasp the apple from a cluttered table with a height of 60 cm.
Figure 17: Task: Cluttered Table. Grasp the apple from a cluttered table with a height of 74.5 cm.
Appendix D Experimental Details
D.1 Evaluation in Simulation
We run large-scale experiments in the simulator to test the spatial generalization curve for each range and determine the minimum data amounts to achieve an optimal success rate. To align with real robot inference, we set the robot’s joints directly to target qpos instead of using environment stepping. We mark the trajectory as success if there is a step both reaching reward $>0.6$ and quat reward $>1.4$. For each policy, we test 48 trajectories and record the final success rate. The experiment consists of three rounds of tests with different seeds, each with 16 trajectories. In each round, the target random points are sampled from a uniform distribution.
D.2 Experiment Result Analysis
We discuss here the analysis of the results of spatial scaling law.
[…]
For the functional relationship between the amount of data required for spatial generalization and spatial volume, we explored two types of fits: a power law and a logarithmic function. The fitting results of power law is given by $y=a\cdot x^{b}=640.32x^{0.35}$ , while the logarithmic fit is $y=a\cdot\ln(b\cdot x)+c=2095.51\cdot\ln{(0.3\cdot x)}-642.93$. We found that the logarithmic fit becomes unreasonable as the spatial volume $x$ approaches 0, where the required data, $y=a\cdot\ln(b\cdot x)+c=2095.51\cdot\ln{(0.3\cdot x)}-642.93$ , tends to negative infinity, which is unrealistic since the amount of data cannot be negative. Furthermore, the fit results in a negative value at $x=3.3$, where $bx=1$ and $\log(bx)=0$, leading to a required data amount of $a\log(bx)+c=-642.93$. In the long run (for larger spatial volumes), the power law fit better describes the relationship between the data required for spatial generalization and spatial volume. Fig. 18 shows the selected power law relationship, and Fig. 19 displays the logarithmic curve.
D.3 Evaluation in Real World
D.3.1 Real-World Success Rate
To verify our experiment results from the simulator, we evaluate our policy in the real robot. For each range, we choose the first optimal policy that achieves 80% success rate in simulator and evaluate with the real robot for 10 randomly chosen points.
Random Placement of Objects in Real-world Evaluation.
To be more specific, we divide the maximum range, namely the full space ( $41cm*30cm*28cm$), into three equal parts in height, two equal parts in width, and three equal parts in length, resulting in a total of $3*2*3=18$ equally sized spatial cuboid ranges at different positions. For each tested spatial range, ten different points are uniformly selected to test the robot’s manipulation success rate.
D.3.2 Real-World Experiment Settings
To verify our methods, we evaluate our policy in the real robot. We evaluate the student policy which is trained on simulation data of full space ( $41cm*30cm*28cm$) with the real robot.
Spatial Generalization.
As illustrated in Fig. 5, we divide the spatial range horizontally into Left (L), Right (R), and Center (C), and vertically into Front (F) and Back (B), and randomly select points within these sections on different heights to test our policy [^1] The table is an elevated table and can therefore be adjusted to any height.. All these points are arbitrarily selected, which means that the target object has an equal chance of appearing at the center or the edge of the table, far from or close to the robot, in positions that require the robotic arm to reach upwards or downwards. We hope this approach reflects the spatial generalization of our policy comprehensively and honestly.
In addition to the verification of spatial generalization, we also test the ability of Object generalization and background generalization. The details of the experiment setting can be found in Tab. 5.
Object Generalization and Background Generalization.
Furthermore, we have chosen objects of various shapes, sizes, and materials as target objects for testing, all of which were previously unseen by our policy. To be more specific, during training, we only encountered apples with sizes of approximately $(6.9cm-7.4cm,6.9cm-7.4cm,6.9cm-7.4cm)$ in the XYZ direction, while the test target objects span a size range of $(5.1cm-7.0cm,5.1cm-9.7cm,6.0cm-22.7cm)$ in the XYZ direction. Additionally, although apples are roughly spherical, the target objects also include cylindrical and pyriform shapes. Moreover, the target objects encompass different materials such as glass and metal. All these objects serve to demonstrate our policy’s capability to generalize to varying shapes, sizes, and materials.
Table 5: Real-World Experiment Settings. Note that we use only the bounding boxes of the objects for both student policy training and inference. Bolded object sizes are never seen in training.
Generalization Type Object Size (cm) of XYZ Setting Description
Object in Simulation Data
Apple (Domain Randomization) (6.9-7.4, 6.9-7.4, 6.9-7.4) Distribute uniformly in a configurable spatial volume
Object Generalization
Yellow Pear (6.5, 5.5, 10.7) h4: Point Random, Minimal setting
Orange (7.0, 6.7, 6.5) h4: Point Random, Minimal setting
Steel Cup (6.5, 9.7, 10.8) h4: Point Random, Minimal setting
Tin Can (5.1,5.1,7.2) h4: Point Random, Minimal setting
Glass Beaker (5.1, 5.1, 6.0) h4: Point Random, Minimal setting
Background Generalization
Apple (7.2, 7.4, 6.9) h4: Point Random, Orange tablecloth
Unseen Apple (7.0, 7.5, 7.0) h4: Point Random, Black tablecloth
Apple (7.2, 7.4, 6.9) h4: Point Random, Playing video
Apple (7.2, 7.4, 6.9) h4: Point Random, Dim light
Apple (7.2, 7.4, 6.9) h4: Point Random, Flittering light
Complex Generalization
Unseen Apple, Unseen Table (7.0, 7.5, 7.0) Point Random, Table 2
Unseen Apple, Unseen Desk (7.0, 7.5, 7.0) Point Random, Multiple objects
Steel Cup, Unseen Desk (6.5, 9.7, 10.8) Point Random, Daily desktop
Plastic Bottle, Unseen Desk (6.4, 6.4, 22.27) Point Random, Daily desktop
The three camera images, as well as the entire trajectory of grasping, are schematized below in Figure 20:
Appendix E Teacher-Student Training Details
For ease of understanding, Alg. 1 is the pseudo algorithm for the ManiBox method.
Algorithm 1 Training and Deployment Algorithm of ManiBox
1:Simulator environment $Sim$, Real robot $Robot$
2:Proximal Policy Optimization (PPO) for training teacher policy
3:Bounding box extractor $BBoxExtractor$
4:Maximum dataset size $N_{max}$, reward function $R$, domain randomization module $DR$
5:
6:Stage 1: Teacher Policy Training
7:Initialize teacher policy $\pi_{\beta}$ with random weights
8:Set $Sim$ environment with reward function $R$ and domain randomization module $DR$
9:while not converged do
10:     Collect trajectory data $\mathcal{D}_{T}$ in $Sim$ using $\pi_{\beta}$
11:     Update $\pi_{\beta}$ using PPO with $\mathcal{D}_{T}$
12:end while
13:
14:Stage 2: Simulation Data Generation
15:Initialize dataset $\mathcal{D}_{S}=\emptyset$
16:while $|\mathcal{D}_{S}|<N_{max}$ do
17:     Reset $Sim$ environment
18:     Apply domain randomization $DR$ to environment
19:     Generate trajectory $\tau=\{(o^{rob}_{t},a_{t},r_{t})\}_{t=1}^{T}$ using $\pi_{\beta}$
20:     Extract bounding boxes $\{o^{vis}_{t}\}_{t=1}^{T}$ using $BBoxExtractor$
21:     Store $\{(o^{vis}_{t},o^{rob}_{t},a_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{S}$
22:end while
23:
24:Stage 3: Student Policy Training
25:Initialize student policy $\pi_{\theta}$ with random weights
26:while not converged do
27:     Sample trajectory $\tau=\{(o^{vis}_{t},o^{rob}_{t},a_{t})\}_{t=1}^{T}$ from $\mathcal{D}_{S}$
28:     Randomly mask some $o^{vis}_{t}$ to simulate detection failures
29:     Update $\pi_{\theta}(o^{vis}_{\leq t},o^{rob}_{\leq t})$ using the masked $\tau$
30:end while
31:
32:Stage 4: Real-World Deployment
33:$t\leftarrow 0$
34:while not terminated do
35:     Extract real-world bounding boxes $o^{vis}_{t}$ using $BBoxExtractor$
36:     Compute action $a_{t}\sim\pi_{\theta}(o^{vis}_{\leq t},o^{rob}_{\leq t})$
37:     Execute action $a_{t}$ on $Robot$
38:     $t\leftarrow t+1$
39:end while
E.1 Teacher Policy Details
@@
We train our teacher policy using PPO with 8,192 parallel environments in Isaac Lab (Mittal et al., 2023). With a carefully designed reward function and applied PPO techniques, the training is completed in around 40 minutes, reaching approximately 2,000 training iterations. The task involves training a robot to grasp an apple randomly placed on a liftable table and transport it to a designated target position (Figure 21).
Figure 21: The simulation environment consists of a room with a table, an apple, and a robot. The task is for the robot to grasp the apple and place it in a specified location.
The advantages of using a teacher policy instead of direct visual RL training can be summarized in three key aspects. First, the teacher policy relies solely on robot and object information as input, significantly reducing the dimensionality of the state space. Second, the incorporation of privileged information and tailored reward functions for critical steps offers more focused guidance during training, thereby narrowing the exploration space. Lastly, visual RL suffers from limited parallelizability, which results in lower overall training efficiency.
E.1.1 Simulation Scenario Setup
@@
The environment is built based on the lift task in Isaac Lab. During training, the scene consists of a table, an apple, and a robot. The robot is initialized at a fixed starting position at the beginning of each episode. The apple’s size, as well as its x,y position, are uniformly randomized across environments, while the height of the table is also uniformly randomized. The target position, however, remains fixed throughout all training episodes. Key steps captured from three camera perspectives are shown in Figures 22, 23, and 24. To support future extensions to more complex tasks, such as obstacle avoidance during grasping, we represent the robot’s behavior with a 30Hz trajectory. Each environment in the simulator is evenly spaced to prevent interference between different training instances. Further details on the randomized ranges can be found in Table 6.
Figure 22: Key steps of grasping in simulation in cam_high view.
Figure 23: Key steps of grasping in simulation in cam_left_wrist view.
Figure 24: Key steps of grasping in simulation in cam_right_wrist view.
@@
E.1.2 Details of Observation and Action
@@ E.1.3 Implementation of PPO
E.1.4 Reward Function Design
To ensure the policy achieves the desired grasping behavior on our robot, we design a reward function that reflects the task’s objectives. As grasping is complicated, it can be divided into a multi-stage reward function: first reaching the object, then closing the jaws, then picking up the object, and finally retracting the robotic arm to a specified position. Definitions of different part of our reward function for the specific object-grasp task are stated as follows:
  • Fingers Open
    $R_{\text{fingers\_open}}=\tanh\left(\frac{\|\mathbf{p}_{\text{object}}-\mathbf{p}_{\text{ee}}\|_{2}}{\sigma}\right)\cdot\sum_{i=1}^{2}\text{finger\_pos}[i]$
    where $\mathbf{p}_{\text{object}}$ represents the position of the object in world coordinates. $\mathbf{p}_{\text{ee}}$ represents the position of the end-effector in world coordinates. finger_pos represents the joint positions of the robot’s fingers. $\sigma=0.1$ is a scaling factor for distance normalization.
  • Reaching Object
    $R_{\text{reaching\_object}}=\left(1-\tanh\left(\frac{\|\mathbf{p}_{\text{object}}-\mathbf{p}_{\text{ee}}\|_{2}}{\sigma}\right)\right)+\left(1-\tanh\left(\frac{\|\mathbf{p}_{\text{object}}-\mathbf{p}_{\text{ee}}\|_{2}}{\sigma/4}\right)\right)$
    where $\mathbf{p}_{\text{object}}$ represents the position of the object in world coordinates. $\mathbf{p}_{\text{ee}}$ represents the position of the end-effector in world coordinates. $\sigma=0.2$ is a scaling factor for distance normalization.
  • Object Goal Tracking
    $R_{\text{goal\_tracking}}=\mathds{1}(\mathbf{p}_{\text{object},z}>z_{\text{init}}+h_{\text{min}})\cdot\left(1-\tanh\left(\frac{\|\mathbf{p}_{\text{goal}}-\mathbf{p}_{\text{object}}\|_{2}}{\sigma}\right)\right)$
    where $\mathbf{p}_{\text{goal}}$ represents the target position of the object in world coordinates. $\mathbf{p}_{\text{object}}$ represents the position of the object in world coordinates. $\mathbf{p}_{\text{object},z}$ represents the $z$-coordinate (height) of the object in world coordinates. $z_{\text{init}}$ is the initial height of the object. $h_{\text{min}}=0.02$ is the minimal height above the initial height required to achieve the goal. $\sigma=0.3$ is a scaling factor for distance normalization.
  • Object Goal Tracking (Fine-Grained)
    $R_{\text{goal\_tracking\_fine}}=\mathds{1}(\mathbf{p}_{\text{object},z}>z_{\text{init}}+h_{\text{min}})\cdot\left(1-\tanh\left(\frac{\|\mathbf{p}_{\text{goal}}-\mathbf{p}_{\text{object}}\|_{2}}{\sigma_{\text{fine}}}\right)\right)$
    where $\mathbf{p}_{\text{goal}}$ represents the target position of the object in world coordinates. $\mathbf{p}_{\text{object}}$ represents the position of the object in world coordinates. $z_{\text{init}}$ is the initial height of the object. $h_{\text{min}}=0.02$ is the minimal height above the initial height required to achieve the goal. $\sigma_{\text{fine}}=0.05$ provides finer scaling for precise tracking.
  • Action Rate Penalty
    $R_{\text{action\_rate}}=-\|\mathbf{a}_{t}-\mathbf{a}_{t-1}\|_{2}^{2}$
    where $\mathbf{a}_{t}$ represents the action at the current time step. $\mathbf{a}_{t-1}$ represents the action at the previous time step.
  • Joint Velocity Penalty
    $R_{\text{joint\_vel}}=-\sum_{i\in\text{joint\_ids}}\dot{q}_{i}^{2}$
    where $\dot{q}_{i}$ represents the joint velocity of the $i$-th joint. joint_ids is the set of joint indices to penalize.
  • Contact Forces
    $R_{\text{contact\_forces}}=\mathds{1}\left(\frac{\text{finger\_dist}}{4}<\text{proj\_len}<\frac{\text{finger\_dist}}{1.25}+0.2\right)\cdot\left(1.2-\tanh\left(\frac{\text{object\_dist}}{0.2}\right)\right)$
    $\cdot\mathds{1}\left(\|\mathbf{p}_{\text{object}}-\mathbf{p}_{\text{top}}\|<0.1\right)\cdot\left(1-\tanh\left(\frac{\|\mathbf{org\_vec}-\mathbf{y}\|}{0.3}\right)\right)$
[…]
  • Close Fingers
    $R_{\text{close\_fingers}}=\mathds{1}(\text{grasp\_force}>0.5)\cdot\left(1-\tanh\left(\frac{\sum_{i=1}^{2}\text{finger\_pos}[i]}{0.1}\right)\right)$
    where grasp_force represents the force exerted by the fingers on the object. finger_pos represents the joint positions of the robot’s fingers.
  • Lift End Effector (EE)
[…]
    where $\mathbf{p}_{\text{ee},z}$ represents the height of the end-effector above the ground.
  • Lift Object
    $R_{\text{lift\_object}}=\mathds{1}(\mathbf{p}_{\text{object},z}>z_{\text{init}}+h_{\text{min}})\cdot\frac{\text{clamp}(\mathbf{p}_{\text{object},z}-z_{\text{init}}-h_{\text{min}},0.0,0.2)}{0.2}$
    where $\mathbf{p}_{\text{object},z}$ represents the height of the object above the ground. $z_{\text{init}}$ is the initial height of the object. $h_{\text{min}}=0.02$ is the minimal height above the initial height required to achieve the goal.
Details of our weight design for the reward function can be found in Table 8
@@ E.1.4 Reward Function Design
E.2 Data Generation Details
@@
To reduce visual interference between parallel environments and minimize the sim-to-real gap, we incorporate a white room during data collection. The room is designed to be spacious enough to accommodate the maximum working space, including robot, table, and apple. Three RGB-D cameras [^2] But we only use RGB information., calibrated with intrinsic and extrinsic parameters matching those of the real robot, capture visual information. During inference, the images from these cameras are processed by YOLO-World to generate object bounding boxes.
@@
E.3 Student Policy Details
@@
E.4 Sim2Real Technique
We employ the student policy for sim-to-real transfer, where the inputs to the policy included proprioception $o^{rob}_{t}$ which consists of joint positions, and the visual input $o^{vis}_{t}$, represented by bounding boxes. The output action is the desired joint position for the next step. Thus, the primary sim-to-real gaps exist in the joint position and bounding box dimensions.
For joint positions, we develop a real-to-sim code that maps the joint positions obtained from teleoperation on a real-world robot to a simulated robot. We evaluate whether the actions performed at specific target positions were consistent between the real-world and simulated environments. For example, when teleoperating to grasp an apple at a precisely measured location, both the real-world and simulated apples are set at the same coordinates. We used the main arm of the real-world robot to observe whether the simulated robot could successfully grasp the apple.
Special handling is required for the gripper, as there are discrepancies between the URDF model of the gripper and its real-world counterpart. In the simulation, the gripper has two joints, while in reality, it functions as a single joint, leading to differences in value interpretation. To address this, we collect 10 data sets for fitting and utilize function interpolation to map the simulated gripper joint positions to those of the real-world gripper. This approach ensures that the state space and action space are aligned between the simulation and the real environment.
The sim-to-real gap for the bounding boxes corresponds to the image sim-to-real gap. We perform hand-eye calibration on the camera to ensure that the cameras in both the simulation and the real-world robot are approximately aligned. For finer adjustments, we select a specific point in front of the robot and a specific-sized object, observing the bounding boxes obtained from both the simulator and the real-world camera. Since we use bounding boxes, other factors such as lighting and background variations in the sim-to-real gap are not a concern. Ultimately, this method of adjustment ensures that the intrinsic and extrinsic parameters of the camera are consistent between the simulation and the real-world environment.
Other methods to ensure successful Sim2Real, such as domain randomization, are described in the previous section (Sec. 3.2).
Appendix F Hardware Details
The robot (cobot_magic) provided by AGILEX is equipped with a dual-arm design, providing a total of 14 degrees of freedom (DoF), with 7 DoF per arm for versatile movement. Each arm has a maximum valid payload of 1500g, and the gripper has a range of up to 80mm, allowing the robot to perform a variety of grasping and manipulation tasks efficiently. The detailed hardware information and an image of the robot can be found in Figure 25 and Table 13. We also thank ARX for their hardware help.
@@ Appendix F Hardware Details
Appendix G Limitations and Discussions
One of the limitations of ManiBox is that the accuracy of the obtained bounding box depends on the capability of the preceding visual model. As the YOLO-World sometimes may not be able to accurately partition the bounding box, ManiBox may grasp the wrong positions due to the wrong input bounding box. Also, the bounding box is mainly for objects with convex shapes and may be insufficient for handling flexible objects and fluids. In the future, a promising direction is to apply more powerful visual models like Grounding DINO and capture more semantic visual signals like segmentation via SAM in ManiBox to improve its ability. Also, for some complicated manipulation tasks like folding up clothes, directly training an RL teacher agent in the simulator may be difficult. Thus how to utilize human demonstration or state machine to get the expert teacher agent is also worth researching.